# FedDyMem: Efficient Federated Learning with Dynamic Memory and Memory-Reduce for Unsupervised Image Anomaly Detection

### v1

@@ III-B Federated UAD
where $\ell(\cdot)$ denotes the local loss function for individual clients. It requires the $n$-th local model $f^{n}(\cdot;{\bf W}^{n})$ to approximate the normal sample distribution, i.e., ${\mathbb{P}_{n}({\cal X}^{train})}$ , under the condition of intra-class bias (refer to 5). For the federated UAD, shown in Fig. 1, the global model’s objective in one-class classification task is to generalize across all clients by constructing (${\mathbb{P}({\{{\cal X}_{n}^{train}}\}_{n=1}^{N})}$. During the testing phase, the result of a sample $x^{test}\in{{\cal D}^{test}}$ can be expressed as follows:
@@
Figure 2: Overview of our proposed FedDyMem framework. During the training phase, a pre-trained feature extractor is employed to extract features from the local training dataset, which are subsequently projected to the target domain using a projection layer. The projected features are processed through a memory generator to construct the local memory bank. In the initialization, the memory bank is further refined and reduced in size by the memory-reduce method. The server collects the local memory banks from multiple clients and aggregates them into a unified global memory bank. During the inference phase, the client utilizes the unified global memory bank to generate anomaly score maps for the test images.
IV Methodology
IV-A Overview
In this article, we propose an efficient federated learning with dynamic memory and memory-reduce for UAD, called FedDyMem. FedDyMem aims to train model parameters on datasets with locally inconsistent feature distributions and generate a global memory bank that produces uniform distributions, thereby achieving high performance on global ${\cal D}^{test}$.
@@ IV-A Overview
On the $n$-th client, as shown in Fig. 2, the feature extractor, projection layer, and memory generator collaboratively generate high-quality memory representations $\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ in the training workflow, where $i$ denotes the $i$-th sample in ${\cal D}_{n}^{train}$ and $t$ denotes the $t$-th rounds. During the initialization phase ($t=0$), all memory features $\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ extracted from training samples are processed through memory-reduce before being send directly to the server. The local memory bank ${\cal M^{\prime}}^{n,0}$ synchronizes with the aggregated memory bank ¯$\bar{\cal M}^{0}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,0}_{reduce}\}_{n=0}^{N}}$ received from the server. After initialization, the extracted memory feature for sample $i$, denoted as ${\cal M}^{n,i,t}$, is utilized in subsequent rounds to compute the local loss $\ell$. This computation is performed in conjunction with the local memory bank ${\cal M^{\prime}}^{n,t-1}$, which is obtained from the server aggregation during the previous round. Following local training, the final updated memory features ${\cal M}^{n,i,t}$ are reduced by the memory-reduce and sent to the server, summarized as ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\to{\cal M}^{n,t}_{reduce}$. On the server, representation sharing [53, 34, 15] provides distinct advantages over parameter aggregation in addressing heterogeneity in unsupervised federated learning, as parameter aggregation methods suffer from privacy and intellectual property (IP) concerns[55, 56, 57] and introduce high communication overhead. Therefore, FedDyMem employs a memory bank sharing to ensure data privacy. In the $t$-th round, the server receives $N$ local memory banks, denoted as $\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}$. As shown in Fig. 2, FedDyMem employs $K$-means clustering algorithm, where $K$ corresponds to the capacity of the local memory, to aggregate the diverse local memory banks into a unified global memory bank, represented as ¯$\bar{\cal M}^{t}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$. The global memory bank ¯$\bar{\cal M}^{t}$ is subsequently distributed to all clients to update their respective local memory banks. Consequently, at any round $t$, for all $n\in N$, the local memory banks are synchronized such that ${\cal M^{\prime}}^{n,t}=\bar{\cal M}^{t}$. During the testing phase, the extracted memory features ${\cal M}^{i}$ are compared against the memory bank $\bar{\cal M}$ using the nearest neighbor search, and the anomaly score ${\cal A}^{i}$ is computed based on the resulting metric, following the anomaly score function in [51, 58].
0:  $n$, $t$,$x_{i}^{n}\in{\cal D}_{n}^{train}$, ${\cal F}(\cdot)$, ${\cal P}^{n}(\cdot,{\bf W}^{\cal P})$, ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$.
0:  ${\cal M}^{i}$.
1:  ${\{{{\cal F}^{(i,l)}}\}}_{l=1}^{L}\leftarrow{\cal F}({{x_{i}^{n}}})$, where $L$ is determined by ${\cal F}$;
2:  ~${\tilde{\cal F}^{i}}\leftarrow$ Compute by Equation (8);
3:  ${\cal P}^{i}\leftarrow$ Compute by Equation (9);
4:  ${\cal M}^{i}\leftarrow{\cal G}^{n}({\cal P}^{i},{\bf W}^{\cal G};{\bf G})$ with Equation (10) (11);
5:  return  ${\cal M}^{i}$
IV-B Methodology for Obtaining ${\cal M}^{n,i,t}$
Locally, FedDyMem employs a frozen CNN model ${\cal F}(\cdot)$ pre-trained on ImageNet to extract multi-level hierarchical features. Motivated by [51], FedDyMem performs multi-level feature fusion on the extracted features. As shown in Fig. 3, for a given sample ${x_{i}^{n}}\in{\mathbb{R}}^{H\times W\times 3}$ which is from the $n$-th client’s ${\cal D}_{n}$, the extracted features are hierarchically organized into $L$ layers (e.g., $L=5$ in the case of ResNet[59]). We define ${\cal F}^{i,l}\in{\mathbb{R}}^{H_{l}\times W_{l}\times C_{l}}$ as the output features of the $l$-th layer, obtained from ${\cal F}({x_{i}^{n}})$, where $i$ denotes the $i$-th sample from local dataset, $l\in L$ denotes the $l$-th layer and $H_{l}$, $W_{l}$, $C_{l}$ are the size and channel of the feature map. In multi-level feature fusion, we employ a straightforward approach combining up-sampling and feature concatenation to extract features from pre-trained CNNs. This process can be expressed as follows:
~${\tilde{\cal F}^{i}}={f_{concat}}(\{{f_{up}}({\cal F}^{i,l},(H_{0},W_{0}))\}_{l=0}^{L}),$ (8)
@@ IV-B Methodology for Obtaining ${\cal M}^{n,i,t}$
Figure 3: The feature extraction stage in FedDyMem. Multi-level features are extracted using a pretrained CNN combined with multi-level feature fusion. These features are then projected into the anomaly detection domain through a projection layer. Subsequently, memory features are generated using a memory generator.
The memory bank-based approach [49, 50, 51] has demonstrated significant performance in UAD. However, the memory banks utilized in these existing methods are non-trainable, meaning they remain unchanged following initialization. Due to the local bias in feature distributions for federated UAD, using the static memory bank will lead to overfitting on domain-specific data, increasing the risk of error. Furthermore, existing methods operate mainly on a discrete feature space, which increases the variations in feature distributions between clients for normal samples [60]. In this article, we propose a memory generator that integrates spatial information and ensures the continuity of feature space. As shown in Fig. 3, the memory generator ${\cal G}(\cdot,{\bf W}^{\cal G};{\bf G})$ utilizes a coordinate convolution layer $f_{coorconv}(\cdot,{\bf W}^{1\times 1})$ to encode spatial location information by incorporating additional coordinate channels, thereby facilitating the network’s ability to learn spatial transformations more effectively[61]. For the projected feature ${\cal P}^{i}$, the result of the coordinate convolution $\hat{{\cal P}^{i}}\in{\mathbb{R}^{H\times W\times C}}$ is computed as $\hat{{\cal P}^{i}}=f_{coorconv}({f_{concat}(\{{{\cal P}^{i},{\bf X},{\bf Y}}\})},{\bf W}^{1\times 1})$, where $\bf X$ and $\bf Y$ represent the Cartesian coordinates of the feature map. Then, to enhance the continuity of the memory bank, the memory generator utilizes a grid-based approach to construct a continuous feature space. We define a trainable grid space as ${\bf G}\in{\mathbb{R}^{H^{\bf G}\times W^{\bf G}\times C}}$, where $H^{\bf G}$ and $W^{\bf G}$ are hyperparameters indicating the size of the continuous space. The memory generator maps $\hat{{\cal P}^{i}}$ to pixel-wise coordinates ${\dot{\cal P}^{i}}\in{\mathbb{R}^{H\times W\times 2}}$ through a mapping function $\phi(\cdot,{\bf W})\colon{\mathbb{R}^{H\times W\times C}}\to{{\mathbb{R}^{H\times W\times 2}}}$, implemented through two $1\times 1$ convolutional layers in our memory generator. A sample function ${s}(\cdot,\cdot)$ is employed to extract features from the continuous space ${\bf G}$ at specific coordinate values ${(p^{i}_{x},p^{i}_{y})}\in{\dot{\cal P}^{i}}$. The normalized coordinates are computed as follows:
@@
[…]
Figure 4: Illustration of Memory-reduce.
IV-C Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
Considering the bias in feature distributions across clients, the training objective is to optimize the local memory generator to produce normal features that are more closely aligned with the shared global memory bank. FedDyMem introduces a simple metric loss to optimize the parameters of both the memory generator and the projection layer, thereby facilitating client models in generating consistent and high-quality memory features. Specifically, the local memory bank for round $t$ is denoted as ${{\cal M^{\prime}}^{n,t-1}}=\{m^{\prime}_{d}\}_{d=0}^{D}\in{\mathbb{R}^{D\times C}}$, where its size is reduced to $D$ after the initialization on the $n$-th client and global aggregation by the server (details of memory-reduce and aggregation are provided in Section IV-D and Section IV-E). Similar to the patch-based memory bank, the memory features ${\cal M}^{n,i,t}$ can be expressed as a collection of patch features, $\{m_{(h,w)}\in{\mathbb{R}^{C}}\}_{h=0,w=0}^{H,W}$. The loss function $\ell$ is defined as follow:
@@ IV-C Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
IV-D Memory-reduce for ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\to{\cal M}^{n,t}_{reduce}$
The earlier memory-based methods [49, 51] involved storing all sample memory features from the training dataset ${\cal D}_{n}^{train}$ within the memory bank. Accordingly, the capacity of the memory bank scales proportionally with the size of the local dataset. For federated UAD, the large-scale dataset significantly challenges communication efficiency. Moreover, the varying sizes of memory banks also introduce challenges in the aggregation process. Motivated by this, FedDyMem incorporates a memory-reduce method to compress the collected memory features from the local dataset. The objective of memory-reduce can be expressed as follows:
@@
0:  $n$, $t$, ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}$, ${{\cal M^{\prime}}^{n,t-1}}$.
0:  ${\cal M}^{n,0}_{reduce}$.
1:  for $i=0$ to ${\left|{\cal D}_{n}^{train}\right|}$ do
2:     if t >0 then
3:        $w_{i,t}\leftarrow$ Compute by Equation (15) with ${\cal M}^{n,i,t}$ and ${{\cal M^{\prime}}^{n,t-1}}$;
4:     else
5:        $w_{i,t}\leftarrow 1$;
6:     end if
7:  end for
8:  ¯${\bar{\cal M}}_{reduce}^{n,i,t}\leftarrow$ Compute by Equation (14) with $\{{w_{i,t}}\}_{i=0}^{N}$ and $\{{\cal M}^{n,i,t}\}_{i=0}^{N}$;
9:  if t >0 then
10:     $\alpha\leftarrow 1/(t+1)$;
11:     ${\cal M}^{n,t}_{reduce}\leftarrow$ Compute by Equation (16) with ¯${\bar{\cal M}}_{reduce}^{n,i,t}$, ${{\cal M^{\prime}}^{n,t-1}}$ and $\alpha$;
12:  else
13:     ${\cal M}^{n,t}_{reduce}\leftarrow{\bar{\cal M}}_{reduce}^{n,i,t}$;
14:  end if
15:  return  ${\cal M}^{n,t}_{reduce}$
To address the memory bank size issue, [58] employs an Exponential Moving Average (EMA) method. However, the reduced memory bank remains sensitive to the sequence in which input samples are processed. In this article, we introduce a dynamic weighted average, called memory-reduce, that optimizes memory size while preserving sample order independence and effectively correlating with the number of communication rounds. As shown in Fig. 4, we aggregate all memory features, denoted as $\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ and apply a dynamic weighted averaging operation to compute the ¯${\bar{\cal M}}^{n,i,t}_{reduce}$ as:
¯${\bar{\cal M}}_{reduce}^{n,i,t}=\sum\limits_{i=0}^{\left|{{{\cal D}}_{n}^{train}}\right|}{{w_{i,t}}{{{\cal M}}^{n,i,t}}}/\sum\limits_{i=0}^{\left|{{{\cal D}}_{n}^{train}}\right|}{{w_{i,t}}},$ (14)
@@
After computing ¯${\bar{\cal M}}_{reduce}^{n,i,t}$, a round-based EMA (rEMA) method is employed to update the memory bank. This approach ensures smoother and more stable updates during the training process. The update rule is defined as:
${\cal M}^{n,t}_{reduce}=\alpha{\bar{\cal M}}_{reduce}^{n,i,t}+(1-\alpha){{\cal M^{\prime}}^{n,t-1}}$ (16)
@@
Algorithm 3 Aggregation on Server $\cal S$ for Round $t$.
0:  $t$ ,${\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$.
0:  ¯$\bar{\cal M}^{t}$.
1:  $m\leftarrow$[];
2:  $H,W,C\leftarrow{\cal M}^{n,t}_{reduce}.size()$;
3:  $K\leftarrow HW$;
4:  for $n=0$ to $N$ do
5:     $\{m_{(h,w)}^{n}\}_{(h=0,w=0)}^{(H,W)}\leftarrow f_{resize}({{\cal M}^{n,t}_{reduce}},(HW,C))$;
6:     $m\leftarrow m\cup\{m_{(h,w)}^{n}\}_{(h=0,w=0)}^{(H,W)}$;
7:  end for
8:  $\{c_{k}\in{\mathbb{R}^{C}}\}_{k=0}^{K}\leftarrow K$-$(m)$;
9:  ¯$\bar{\cal M}^{t}\leftarrow{f_{resize}}(\{c_{k}\}_{k=0}^{K},(H,W,C))$;
10:  return  ¯$\bar{\cal M}^{t}$
Algorithm 4 Initialization at Round 00.
0:  ${\{{\cal D}_{n}^{train}\}}_{n=0}^{N}$, ${\cal F}(\cdot)$, ${\cal P}^{n}(\cdot,{\bf W}^{\cal P})$, ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$.
0:  ${\{{\cal M^{\prime}}^{n,0}\}}_{n=0}^{N}$
1:  for each client n in parallel do
2:     Initialize ${\bf W}^{\cal P}$, ${\bf W}^{\cal G}$ randomly for ${\cal P}^{n}(\cdot,{\bf W}^{\cal P})$, ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$;
3:     Initialize $\bf G$ by Xavier normal initialization method[62] for ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$;
4:     for $i=0$ to $\left|{\cal D}_{n}^{train}\right|$ do
5:        ${\cal M}^{i}\leftarrow$ Compute by Algorithm 1 with $t=0$, $n$, $x_{i}^{n}$, ${\cal F}(\cdot)$, ${\bf W}^{\cal P}$, ${\bf W}^{\cal G}$, ${\bf G}$;

### v2

@@ 3.2 Federated UAD
where $\ell(\cdot)$ denotes the local loss function for individual clients. It requires the $n$-th local model $f^{n}(\cdot;{\bf W}^{n})$ to approximate the normal sample distribution, i.e., $\mathbb{P}_{n}({\cal X}^{train})$} , under the condition of intra-class bias (refer to 5). For the federated UAD, shown in Fig. 1, the global model’s objective in one-class classification task is to generalize across all clients by constructing $\mathbb{P}({\{{\cal X}_{n}^{train}}\}_{n=1}^{N})$}. During the testing phase, the result of a sample $x^{test}\in{{\cal D}^{test}}$ can be expressed as follows:
@@
4 Methodology
4.1 Overview
@@
In this article, we propose an efficient federated learning with dynamic memory and memory-reduce for UAD, called FedDyMem. FedDyMem aims to train model parameters on datasets with locally inconsistent feature distributions and generate a global memory bank that produces uniform distributions, thereby achieving high performance on global ${\cal D}^{test}$. Each communication round in FedDyMem comprises the following three steps of training: 1) Local training: On the $n$-th client, as shown in Fig. 2, the feature extractor, projection layer, and memory generator collaboratively generate high-quality memory representations $\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ in the training workflow, where $i$ denotes the $i$-th sample in ${\cal D}_{n}^{train}$ and $t$ denotes the $t$-th rounds. During the initialization phase ($t=0$), all memory features $\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ extracted from training samples are processed through memory-reduce before being send directly to the server. The local memory bank ${\cal M^{\prime}}^{n,0}$ synchronizes with the aggregated memory bank $\bar{\cal M}^{0}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,0}_{reduce}\}_{n=0}^{N}}$ received from the server. After initialization, the extracted memory feature for sample $i$, denoted as ${\cal M}^{n,i,t}$, is utilized in subsequent rounds to compute the local loss $\ell$. This computation is performed in conjunction with the local memory bank ${\cal M^{\prime}}^{n,t-1}$, which is obtained from the server aggregation during the previous round. 2) Memory preparation: After the local training step in each communication round, both the projection layer and the memory generator are updated. Subsequently, the local training samples are passed through the trained models to extract the corresponding memory features, denoted as ${\cal M}^{n,i,t}$. These per-sample memory features are then reduced by the memory-reduce, summarized as ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\to{\cal M}^{n,t}_{reduce}$. Representation sharing ([kim2023protofl, tan2022fedproto, zhang2024prototype]) provides distinct advantages over parameter aggregation in addressing heterogeneity in unsupervised federated learning, as parameter aggregation methods suffer from privacy and intellectual property (IP) concerns ([li2021survey, zhang2024fedtgp]) and introduce high communication overhead. Therefore, FedDyMem introduces the memory-reduce operation and the memory bank sharing mechanism to mitigate communication costs while preserving client data privacy. 3) Aggregation and distribution: In the $t$-th round, the server receives $N$ local memory banks, denoted as $\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}$. As shown in Fig. 2, FedDyMem employs $K$-means clustering algorithm, where $K$ corresponds to the capacity of the local memory, to aggregate the diverse local memory banks into a unified global memory bank, represented as $\bar{\cal M}^{t}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$. The global memory bank $\bar{\cal M}^{t}$ is subsequently distributed to all clients to update their respective local memory banks. Consequently, at any round $t$, for all $n\in N$, the local memory banks are synchronized such that ${\cal M^{\prime}}^{n,t}=\bar{\cal M}^{t}$.
During the testing phase, the extracted memory features ${\cal M}^{i}$ are compared against the memory bank $\bar{\cal M}$ using the nearest neighbor search, and the anomaly score ${\cal A}^{i}$ is computed based on the resulting metric, following the anomaly score function in ([roth2022towards, lee2022cfa]).
Figure 3: Illustration of the feature extraction process in FedDyMem.
4.2 Methodology for Obtaining ${\cal M}^{n,i,t}$
[…]
${\tilde{\cal F}^{i}}={f_{concat}}(\{{f_{up}}({\cal F}^{i,l},(H_{0},W_{0}))\}_{l=0}^{L}),$ (8)
@@ 4.2 Methodology for Obtaining ${\cal M}^{n,i,t}$
[…]
7:return ${\cal M}^{i}$
The memory bank-based approach ([cohen2020sub, defard2021padim, roth2022towards]) has demonstrated significant performance in UAD. However, the memory banks utilized in these existing methods are non-trainable, meaning they remain unchanged following initialization. Due to the local bias in feature distributions for federated UAD, using the static memory bank will lead to overfitting on domain-specific data, increasing the risk of error. Furthermore, existing methods operate mainly on a discrete feature space, which increases the variations in feature distributions between clients for normal samples ([lee2025continuous]). In this article, we propose a memory generator that integrates spatial information and ensures the continuity of feature space. As shown in Fig. 3, the memory generator ${\cal G}(\cdot,{\bf W}^{\cal G};{\bf G})$ utilizes a coordinate convolution layer $f_{coorconv}(\cdot,{\bf W}^{1\times 1})$ to encode spatial location information by incorporating additional coordinate channels, thereby facilitating the network’s ability to learn spatial transformations more effectively ([liu2018intriguing]). For the projected feature ${\cal P}^{i}$, the result of the coordinate convolution $\hat{{\cal P}^{i}}\in{\mathbb{R}^{H\times W\times C}}$ is computed as $\hat{{\cal P}^{i}}=f_{coorconv}({f_{concat}(\{{{\cal P}^{i},{\bf X},{\bf Y}}\})},{\bf W}^{1\times 1})$, where $\bf X$ and $\bf Y$ represent the Cartesian coordinates of the feature map. Then, to enhance the continuity of the memory bank, the memory generator utilizes a grid-based approach to construct a continuous feature space. We define a trainable grid space as ${\bf G}\in{\mathbb{R}^{H^{\bf G}\times W^{\bf G}\times C}}$, where $H^{\bf G}$ and $W^{\bf G}$ are hyperparameters indicating the size of the continuous space. The memory generator maps $\hat{{\cal P}^{i}}$ to pixel-wise coordinates ${\dot{\cal P}^{i}}\in{\mathbb{R}^{H\times W\times 2}}$ through a mapping function $\phi(\cdot,{\bf W})\colon{\mathbb{R}^{H\times W\times C}}\to{{\mathbb{R}^{H\times W\times 2}}}$, implemented through two $1\times 1$ convolutional layers in our memory generator. A sample function ${s}(\cdot,\cdot)$ is employed to extract features from the continuous space $\bf G$} at specific coordinate values ${(p^{i}_{x},p^{i}_{y})}\in{\dot{\cal P}^{i}}$. The normalized coordinates are computed as follows:
@@
where $\left\lfloor{\cdot}\right\rfloor$ represents the floor operation. Finally, the features ${\cal O}^{i}$, sampled from the continuous feature space, are concatenated with the coordinate convolution outputs $\hat{{\cal P}}^{i}$. This concatenated tensor is then passed through a convolutional layer to generate the memory feature ${\cal M}^{i}\in\mathbb{R}^{H\times W\times C}$ for client $n$ at round $t$. The complete procedure for obtaining the memory feature is presented in Algorithm 1.
4.3 Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
Considering the bias in feature distributions across clients, the training objective is to optimize the local memory generator to produce normal features that are more closely aligned with the shared global memory bank. FedDyMem introduces a simple metric loss to optimize the parameters of both the memory generator and the projection layer, thereby facilitating client models in generating consistent and high-quality memory features. Specifically, the local memory bank for round $t$ is denoted as ${{\cal M^{\prime}}^{n,t-1}}=\{m^{\prime}_{d}\}_{d=0}^{H\times W}\in{\mathbb{R}^{{(H\times W)}\times C}}$, where its size is reduced to $H\times W$ after the initialization on the $n$-th client and global aggregation by the server (details of memory-reduce and aggregation are provided in Section 4.4 and Section 4.5). Similar to the patch-based memory bank, the memory features ${\cal M}^{n,i,t}$ can be expressed as a collection of patch features, $\{m_{(h,w)}\in{\mathbb{R}^{C}}\}_{h=0,w=0}^{H,W}$. The loss function $\ell$ is defined as follow:
@@ 4.3 Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
4.4 Memory-reduce for ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\to{\cal M}^{n,t}_{reduce}$
The earlier memory-based methods ([cohen2020sub, roth2022towards]) involved storing all sample memory features from the training dataset ${\cal D}_{n}^{train}$ within the memory bank. Accordingly, the capacity of the memory bank scales proportionally with the size of the local dataset. For federated UAD, this design leads to high communication overhead when applied to large-scale datasets and also introduces privacy risks. When each memory feature corresponds directly to a raw input sample, it becomes vulnerable to privacy leakage under reconstruction or inversion attacks. Moreover, the varying sizes of memory banks also introduce challenges in the aggregation process. Motivated by this, FedDyMem incorporates a memory-reduce method to compress the collected memory features from the local dataset, thereby reducing the mutual information between the memory features and the raw data. The objective of memory-reduce can be expressed as follows:
@@
Figure 4: Illustration of Memory-reduce.
To address the memory bank size issue, ([lee2022cfa]) employs an Exponential Moving Average (EMA) method. However, the reduced memory bank remains sensitive to the sequence in which input samples are processed. In this article, we introduce a dynamic weighted average, called memory-reduce, that optimizes memory size while preserving sample order independence and effectively correlating with the number of communication rounds. As shown in Fig. 4, we aggregate all memory features, denoted as $\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$ and apply a dynamic weighted averaging operation to compute the ${\bar{\cal M}}^{n,t}_{reduce}$ as:
${\bar{\cal M}}_{reduce}^{n,t}=\sum\limits_{i=0}^{\left|{{{\cal D}}_{n}^{train}}\right|}{{w_{i,t}}{{{\cal M}}^{n,i,t}}}/\sum\limits_{i=0}^{\left|{{{\cal D}}_{n}^{train}}\right|}{{w_{i,t}}},$ (14)
@@
1:$n$, $t$, $\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$}, ${\cal M^{\prime}}^{n,t-1}$}.
[…]
3:for $i=0$ to $\left|{\cal D}_{n}^{train}\right|$} do
4:  if t >0 then
5:   $w_{i,t}\leftarrow$ Compute by Equation (15) with ${\cal M}^{n,i,t}$ and ${\cal M^{\prime}}^{n,t-1}$};
6:  else
7:   $w_{i,t}\leftarrow 1$;
8:  end if
9:end for
[…]
11:if t >0 then
12:  $\alpha\leftarrow 1/(t+1)$;
13:  ${\cal M}^{n,t}_{reduce}\leftarrow$ Compute by Equation (16) with ${\bar{\cal M}}_{reduce}^{n,t}$, ${\cal M^{\prime}}^{n,t-1}$} and $\alpha$;
14:else
15:  ${\cal M}^{n,t}_{reduce}\leftarrow{\bar{\cal M}}_{reduce}^{n,t}$;
16:end if
17:return ${\cal M}^{n,t}_{reduce}$
Algorithm 3 Aggregation on Server $\cal S$ for Round $t$.
1:$t$ $\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}$}.
[…]
6:for $n=0$ to NN do
7:  $\{m_{(h,w)}^{n}\}_{(h=0,w=0)}^{(H,W)}\leftarrow f_{resize}({{\cal M}^{n,t}_{reduce}},(HW,C))$;
8:  $m\leftarrow m\cup\{m_{(h,w)}^{n}\}_{(h=0,w=0)}^{(H,W)}$;
9:end for
[…]
12:return $\bar{\cal M}^{t}$
After computing ${\bar{\cal M}}_{reduce}^{n,t}$, a round-based EMA (rEMA) method is employed to update the memory bank. This approach ensures smoother and more stable updates during the training process. The update rule is defined as:
${\cal M}^{n,t}_{reduce}=\alpha{\bar{\cal M}}_{reduce}^{n,t}+(1-\alpha){{\cal M^{\prime}}^{n,t-1}},$ (16)
@@
Algorithm 4 Initialization at Round 0.
[…]
3:for each client n in parallel do
4:  Initialize ${\bf W}^{\cal P}$, ${\bf W}^{\cal G}$ randomly for ${\cal P}^{n}(\cdot,{\bf W}^{\cal P})$, ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$;
5:  Initialize $\bf G$ by Xavier normal initialization for ${\cal G}^{n}(\cdot,{\bf W}^{\cal G};{\bf G})$;
6:  for $i=0$ to $\left|{\cal D}_{n}^{train}\right|$ do
[…]
8:  end for
9:  ${\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\leftarrow[{\cal M}^{0},{\cal M}^{1},\cdots]$;
10:  ${\cal M}^{n,0}_{reduce}\leftarrow$ Compute by Algorithm 2 with $n$, $t=0$ and $\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}$};
11:  ${\bf W}^{{\cal P},n,0}\leftarrow{\bf W}^{\cal P}$, ${\bf W}^{{\cal G},n,0}\leftarrow{\bf W}^{\cal G}$, ${{\bf G}^{n,0}}\leftarrow{\bf G}$
12:end for
[…]
15:for $n=0$ to NN do
16:  ${\cal M^{\prime}}^{n,0}\leftarrow\bar{\cal M}^{0}$;
17:end for
18:return ${\{{\cal M^{\prime}}^{n,0}\}}_{n=0}^{N}\leftarrow[{\cal M^{\prime}}^{0,0},{\cal M^{\prime}}^{1,0},\cdots]$
4.5 $\bar{\cal M}^{t}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$ on Server
At round $t$, the server $\cal S$ collects the newly generated memory banks from $N$ clients, denoted as ${\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}\in{{\mathbb{R}^{N\times H\times W\times C}}}$. The objective of the server is to aggregate these memory banks into a unified, representative memory bank $\bar{\cal M}^{t}$. This process is formalized as ${\cal S}(\cdot,\cdot)\colon{{\mathbb{R}}^{N\times H\times W\times C}}\to{{\mathbb{R}}^{H\times W\times C}}$. Representation-based federated learning approaches, such as those proposed in ([tan2022fedproto, tan2022federated]), utilize weighted aggregation of class prototypes during model updates. However, significant feature bias in the early training stages leads to aggregated memory that deviates significantly from the client-specific memory, which often contains incomplete product types. This misalignment hinders convergence, posing a challenge for clients with limited data diversity. Thus, we employ a clustering-based approach to aggregate the collected memory banks effectively. Specifically, the memory features are represented as $N\times H\times W$ patch features, denoted by ${m_{(h,w)}^{n}}\in{\mathbb{R}^{C}}$. These patch features are subsequently clustered using the $K$-means algorithm, where the number of clusters, $K$, corresponds to $H\times W$. The clustering results indicate the overall distribution and similarity relationships between the patch-level features stored across all banks at the current stage of the process. Cluster centers describe the representative features of a memory bank. Let $c_{k}\in{\mathbb{R}^{C}}$ denote the center of the $k$-th cluster, serving as a global memory bank updated by $\bar{\cal M}^{t}={f_{resize}}(\{c_{k}\}_{k=0}^{K},(H,W,C))$ and shared across individual clients, $\forall n\in N,\ {\cal M^{\prime}}^{n,t}=\bar{\cal M}^{t}$. The aggregation methodology is outlined in Algorithm 3, while the comprehensive training procedure for FedDyMem is presented in Algorithm 5.
Algorithm 5 Training for FedDyMem.
1:NN Clients $\{{\cal C}_{n}\}_{n=1}^{N}$ with ${\{{\cal D}_{n}^{train}\}}_{n=0}^{N}$, one server $\cal S$, the number of rounds $T$, the local epochs $E$.
[…]
3:for $t=1$ to $T$ do
4:  for each client $n$ in parallel do
5:   ${\cal M}^{n,t}_{reduce}\leftarrow$ ClientUpdate(n, t, E, ${\cal D}_{n}^{train}$);